OpenStudy (anonymous):

A rectangle has a length of 18 cm and a perimeter of 68 cm. Which equation could be used to find the width of the rectangle? A. w + 18 = 68 B. 18w = 68 C. 2w + 18 = 68 D. 2w + 36 = 68

OpenStudy (anonymous):

A rectangle has a length of 18 cm and a perimeter of 68 cm. Which equation could be used to find the width of the rectangle? A. w + 18 = 68 B. 18w = 68 C. 2w + 18 = 68 D. 2w + 36 = 68

OpenStudy (anonymous):

I would say choice A!

OpenStudy (anonymous):

perimeter is sum of width and length, so obviously not double it at all and 36 is obvious to eliminate!

OpenStudy (anonymous):

Medal? :D

OpenStudy (anonymous):

thats wrong